# Transcendence and normality of complex numbers via Hurwitz continued fractions

### v2

@@ Abstract.
We study the topological, dynamical, and descriptive set theoretic properties of Hurwitz continued fractions. Hurwitz continued fractions associate an infinite sequence of Gaussian integers to every complex number which is not a Gaussian rational. The resulting space of sequences of Gaussian integers $\Omega$ is not closed. By means of an algorithm, we show that $\Omega$ contains a natural subset whose closure $\overline{\mathsf{R}}$ encodes continued fraction expansions of complex numbers which are not Gaussian rationals. We prove that $(\overline{\mathsf{R}},\sigma)$ is a subshift with a feeble specification property. As an application, we determine the rank in the Borel hierarchy of the set of Hurwitz normal numbers with respect to the complex Gauss measure. We also construct a family of complex transcendental numbers with bounded partial quotients.
@@ 1. Introduction
where all the terms $a_{n}$ are integers and $a_{n}\geq 1$ for any $n\geq 1$. Regular continued fractions are a powerful tool in number theory. They provide a number system where the expansion of a real number is infinite if and only if it is irrational, and eventually periodic if and only if the number is quadratic; for example,
@@
There have been several attempts to obtain multidimensional continued fraction algorithms (see, for example, [24, 36]) but the higher dimensional theory, in comparison, is hardly as comprehensive as the one-dimensional theory. One such example is the continued fraction theory for complex plane that we focus on in this paper. In his 1887 paper [21], Hurwitz introduced a complex continued fraction algorithm where all the partial quotients $(a_{n})_{n\geq 1}$ are Gaussian integers. Despite its age, Hurwitz algorithm had not received extensive attention until recently, when interesting discoveries were made regarding its Diophantine approximation and dynamical properties [9, 12, 13, 16, 20]. Hurwitz showed that his continued fractions associate to each number $z\in\mathbb{C}\setminus\mathbb{Q}(i)$ a unique sequence $(a_{n})_{n\geq 0}$ in $\mathbb{Z}[i]$ satisfying
@@
Hurwitz continued fractions can also be approached from a dynamical perspective. In fact, the role of the unit interval is played by the square
@@
and the definition of the Hurwitz map $T$ is similar to that of $T_{\mathbb{R}}$ (see Section 3). We say a sequence $(a_{n})_{n\geq 1}$ in $\mathbb{Z}[i]$ is valid when $z=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$ for some $z\in\mathfrak{F}\setminus\mathbb{Q}(i)$ and we denote by $\Omega$ the space of valid sequences. A central difference between regular and Hurwitz continued fractions is that, while $\mathbb{N}^{\mathbb{N}}$ is the space of valid sequences for regular continued fractions, the set $\Omega$ is not even closed in $\mathbb{Z}[i]^{\mathbb{N}}$ (see Section 4). Working with closed sets is often crucial for developing a natural dynamical perspective. One solution might involve taking the closure of $\Omega$, but this approach can result in several intricate points that prove challenging to handle. Nevertheless, it turns out that there is a naturally closed subset related to $\Omega$ that still encapsulates its fundamental characteristics. To be more precise, for each $m\in\mathbb{N}$ and $(b_{1},\ldots,b_{m})\in\mathbb{Z}[i]^{m}$, define the cylinder of level $m$ based on $\mathbf{b}$ as the set
@@
$\overline{\Lambda}\left(\overline{\mathsf{R}}\right)=\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$.
The containment $\overline{\Lambda}\left(\overline{\mathsf{R}}\right)\subseteq\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$ will be a consequence of the approximation properties of Hurwitz continued fractions (Proposition 3.1.v). We will prove $\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)\subseteq\overline{\Lambda}\left(\overline{\mathsf{R}}\right)$, using an algorithm (Algorithm 4.3) whose input is a sequence $\mathbf{a}\in\Omega\setminus\overline{\mathsf{R}}$ and whose output is a sequence $\mathbf{b}\in\overline{\mathsf{R}}$ such that $\overline{\Lambda}(\mathbf{b})=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$. In [9, Corollary 6.4], $\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)\subseteq\overline{\Lambda}\left(\overline{\mathsf{R}}\right)$ was obtained by different means but the argument there fails to provide for any $\mathbf{a}=(a_{n})_{n\geq 1}\in\Omega\setminus\overline{\mathsf{R}}$ a sequence $\mathbf{b}\in\overline{\mathsf{R}}$ for which $\overline{\Lambda}(\mathbf{b})=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$.
Let $\sigma\colon\mathbb{Z}[i]^{\mathbb{N}}\to\mathbb{Z}[i]^{\mathbb{N}}$ be the left-shift map (the map that erases the first Gaussian integer). In [38], Yuri proved that $(\overline{\mathsf{R}},\sigma)$ is a sofic shift realized by a directed graph whose vertices are open prototype sets (see definition in Section 3) and with edges labelled by Gaussian integers. Yuri’s paper is concerned with the ergodic theory of a class of dynamical systems defined in compact subsets of $\mathbb{R}^{n}$, $n\in\mathbb{N}$. She considered measures equivalent to the $n$-dimensional Lebesgue measure and, thus, certain Lebesgue-null sets as well as some subsets of the associated shift space are discarded. These omissions force us to make a more detailed study of the relationship between $(\overline{\mathsf{R}},\sigma)$ and the dynamics of Hurwitz continued fraction when dealing with other problems such as those in Section 4 and Section 6.
@@
In fact, we will establish a more robust result by categorizing the set of normal numbers within the framework of the Borel hierarchy. The Borel hierarchy serves as a natural rank, quantifying the number of quantifiers required to construct a Borel set, commencing from open sets. In some sense, this measures how difficult it is to determine whether a complex number is a Hurwitz normal number. A related result was obtained by Dani and Nogueira when they were studying the image of $\mathbb{Z}[i]^{2}$ under binary quadratic forms [13, Propositions 8.5, 8.6].
@@
The subshift $(\overline{\mathsf{R}},\sigma)$ is not a full shift, in other words, $\overline{\mathsf{R}}$ is not of the form $\mathscr{A}^{\mathbb{N}}$ for some $\mathscr{A}\subseteq\mathbb{Z}[i]$. Nonetheless, Theorem 1.3 shows that the system $(\overline{\mathsf{R}},\sigma)$ still has a considerable level of freedom in terms of specification. The specification property was introduced by Bowen in his seminal paper [6]. Bowen’s Specification Theorem asserts that all transitive hyperbolic systems possess the “specification property”, wherein complete independence is established among finite subsets with bounded gaps. The specification and its variations have proven to be potent tools for unlocking access to various other properties, such as the uniqueness of measures of maximal entropy [6, 11], universality [32], and the intricate structure of generic points [4, 14].
The second application of Theorem 1.1 is related to transcendental numbers. In [2], Adamczewski and Bugeaud solved Cobham’s Conjecture when they proved that, for any $b\in\mathbb{N}_{\geq 2}$, a real number whose expansion in base $b$ is non-periodic and generated by finite automaton must be transcendental. They extended their work to continued fractions [1] and, afterwards in [8], Bugeaud gave a combinatorial condition on non-periodic continued fraction ensuring its transcendence. This condition can be expressed in an extremely neat manner using the repetition exponent introduced by Bugeaud and Kim [10]. Given a finite alphabet $\mathscr{A}$ and a sequence $\mathbf{a}=(a_{n})_{n\geq 1}$ in $\mathscr{A}$, for each $n\in\mathbb{N}$ define
@@
Bugeaud proved that for any bounded non-periodic sequence $\mathbf{a}=(a_{n})_{n\geq 1}\in\mathbb{N}^{\mathbb{N}}$, if $\operatorname{rep}(\mathbf{a})<\infty$, the continued fraction (1) is transcendental for any $a_{0}\in\mathbb{Z}$ [8, Theorem 1.3]. In [19], the second author obtained a complex version of this result under the addition assumption $\min_{n}|a_{n}|\geq\sqrt{8}$. In this paper, we apply the algorithm from the proof of Theorem 1.1 to construct a family of transcendental numbers whose partial quotients $(a_{n})_{n\geq 1}$ are bounded and such that $|a_{n}|=2$ for infinitely many $n\in\mathbb{N}$. Schmidt’s Subspace Theorem, a deep result in Diophantine approximation, lies at the heart of this construction.
@@
The structure of the paper is as follows. First, in Section 2, we establish notation and conventions. In Section 3, we introduce the basic definitions of Hurwitz continued fractions, we discuss the classification of cylinders and prototype sets, and we give some basic results. In Section 4 we prove Theorem 1.1 and state Algorithm 4.3 the algorithm without proof. Section 5 starts with some basic notions on Borel Hierarchy and the feeble-specification property. Later, we show Theorem 1.3 and then Theorem 5.1 (which implies Theorem 1.2). In Section 6 we show Theorem 1.4. Lastly, Section 7 contains the proof Algorithm 4.3.
@@ 2. Notation and conventions
    Given finitely many functions $h_{1},\ldots,h_{n}$ that are either $\iota$, $\tau_{a}$ for some $a\in\mathbb{Z}[i]$ or some element in $\mathsf{Di}_{8}$ (see Section 3), we write $h_{n}\cdots h_{1}:=h_{n}\circ\cdots\circ h_{1}$.
@@
  8. viii.
    If $X$ is a topological space and $A\subseteq X$, we let $\operatorname{Cl}_{X}(A)$ (resp. $\operatorname{Int}_{X}(A)$ ) be the closure of $A$ (resp. the interior). We omit the sub-index $X$ if the space is clear.
@@
    The term word typically refers to a finite set of ordered symbols, while a sequence denotes an infinite one. However, in certain situations where the set may be finite or infinite, we use both terms interchangeably.
@@
We use implicitly the formulas for the inversion of discs and lines in the complex plane throughout the paper. Namely, for $k\in\mathbb{R}$, we have
@@ 3. Hurwitz continued fractions
$\quad T(z)=\begin{cases}z^{-1}-[z^{-1}],&\text{ if }z\neq 0,\\ 0,&\text{ if }z=0.\end{cases}$
For any $z\in\mathfrak{F}\setminus\{0\}$, we define $a_{0}(z)=0$, $a_{1}(z)\colon=[z^{-1}]$ and, if $T^{n-1}(z)\neq 0$, we put $a_{n}(z)\colon=a_{1}(T^{n-1}(z))$ for $n\in\mathbb{N}$. As it is customary, the exponent in $T$ denotes iteration. For a general $z\in\mathbb{C}$, we define $a_{0}(z)=[z]$ and $a_{n}(z)=a_{n}(z-[z])$, $n\in\mathbb{N}$. The Hurwitz continued fraction (HCF) of $z$ is
$[a_{0}(z);a_{1}(z),a_{2}(z),\ldots]_{\mathbb{C}}\colon=\cfrac{1}{a_{1}(z)+\cfrac{1}{a_{2}(z)+\cfrac{1}{\ddots}}}.$
We will restrict our attention to $z\in\mathfrak{F}$. Note that
@@
For such $\xi$, we have defined $[\xi]$ to be the solution of (7) with the largest real and imaginary parts. In many situations, the breaking ties criterion is irrelevant for it applies on a set of Hausdorff dimension $1$, hence Lebesgue null. However, it ensures the uniqueness of the HCF expansion. If we are willing to loosen slightly the uniqueness of the expansion, Theorem 1.1 and its proof imply that it makes no difference what solution of (7) we choose.
We will use continued fractions which are not necessarily HCFs and denote them by
@@
We use the notation $[b_{0};b_{1},b_{2},\ldots]_{\mathbb{C}}$ for HCFs.
@@
Let $z=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}\in\mathfrak{F}\setminus\mathbb{Q}(i)$ and let $(p_{n})_{n\geq 0}$, $(q_{n})_{n\geq 0}$ be as in (5).
  1. i.
@@
  4. iv.
    [13, Corollary 3.3] We have
    $\forall n\in\mathbb{N}\quad z=\frac{(a_{n+1}+[0;a_{n+2},a_{n+3},\ldots])p_{n}+p_{n-1}}{(a_{n+1}+[0;a_{n+2},a_{n+3},\ldots])q_{n}+q_{n-1}}.$
  5. v.
    [29, Theorem 1] If $z=[0;a_{1},a_{2},\ldots]\in\mathfrak{F}\setminus\mathbb{Q}(i)$, then
@@
    $\forall n\in\mathbb{N}\quad\left|z-\frac{p_{n}}{q_{n}}\right|<\frac{\lambda}{|q_{n}q_{n+1}|^{2}}.$
  7. vii.

### v3

@@ Abstract.
We study the topological, dynamical, and descriptive set-theoretic properties of Hurwitz continued fractions. Hurwitz continued fractions associate an infinite sequence of Gaussian integers to every complex number which is not a Gaussian rational. The resulting space of sequences of Gaussian integers $\Omega$ is not closed. Using an iterative procedure, we show that $\Omega$ contains a natural subset whose closure $\overline{\mathsf{R}}$ encodes continued fraction expansions of complex numbers which are not Gaussian rationals. We prove that $(\overline{\mathsf{R}},\sigma)$ is a subshift with a feeble specification property. As an application, we determine the rank in the Borel hierarchy of the set of Hurwitz normal numbers with respect to the complex Gauss measure. We also construct a family of complex transcendental numbers with bounded partial quotients.
@@ 1. Introduction
where all the terms $a_{n}$ are integers and $a_{n}\geq 1$ for any $n\geq 1$. Regular continued fractions are a powerful tool in number theory. They provide a number system where the expansion of a real number is infinite if and only if it is irrational, and eventually periodic if and only if the number is a quadratic irrational; for example,
@@
There have been several attempts to obtain multidimensional continued fraction algorithms (see, for example, [24, 36]). Still, the higher-dimensional theory, in comparison, is hardly as comprehensive as the one-dimensional theory. One such example is the continued fraction theory for the complex plane we focus on. In his 1887 paper [21], Hurwitz introduced a complex continued fraction algorithm where all the partial quotients $(a_{n})_{n\geq 1}$ are Gaussian integers. Despite its age, the Hurwitz algorithm had not received extensive attention until recently, when interesting discoveries were made regarding its Diophantine approximation and dynamical properties [9, 12, 13, 16, 20]. Hurwitz showed that his continued fractions associate to each number $z\in\mathbb{C}\setminus\mathbb{Q}(i)$ a unique sequence $(a_{n})_{n\geq 0}$ in $\mathbb{Z}[i]$ satisfying
@@
Like regular continued fractions, Hurwitz continued fractions can also be approached dynamically. The square
@@
plays the role of the unit interval, and the definition of the Hurwitz map $T$ is similar to that of $T_{\mathbb{R}}$ (see Section 3). We say a sequence $(a_{n})_{n\geq 1}$ in $\mathbb{Z}[i]$ is valid when $z=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$ for some $z\in\mathfrak{F}\setminus\mathbb{Q}(i)$ and we denote by $\Omega$ the space of valid sequences. A central difference between regular and Hurwitz continued fractions is that, while $\mathbb{N}^{\mathbb{N}}$ is the space of valid sequences for regular continued fractions, the set $\Omega$ is not even closed in $\mathbb{Z}[i]^{\mathbb{N}}$ (see Section 4). Working with closed sets is often crucial for developing a natural dynamical perspective. One solution might involve taking the closure of $\Omega$, but this approach can result in several intricate points that prove challenging to handle. Nevertheless, it turns out that there is a naturally closed subset related to $\Omega$ that still encapsulates its fundamental characteristics. To be more precise, for each $m\in\mathbb{N}$ and $\mathbf{b}=(b_{1},\ldots,b_{m})\in\mathbb{Z}[i]^{m}$, define the cylinder of level $m$ based on $\mathbf{b}$ as the set
@@
$\overline{\Lambda}\left[\,\overline{\mathsf{R}}\,\right]=\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$.
The inclusion $\overline{\Lambda}\left[\,\overline{\mathsf{R}}\,\right]\subseteq\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)$ will be a consequence of the approximation properties of Hurwitz continued fractions (Proposition 3.1.(iv)). We will prove $\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)\subseteq\overline{\Lambda}\left[\,\overline{\mathsf{R}}\,\right]$, using an iterative procedure (Procedure 4.3) whose input is a sequence $\mathbf{a}\in\Omega\setminus\overline{\mathsf{R}}$ and whose output is a sequence $\mathbf{b}\in\overline{\mathsf{R}}$ such that $\overline{\Lambda}(\mathbf{b})=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$. In [9, Corollary 6.4], $\overline{\mathfrak{F}}\setminus\mathbb{Q}(i)\subseteq\overline{\Lambda}\left[\,\overline{\mathsf{R}}\,\right]$ was obtained by different means but the argument there fails to provide for any $\mathbf{a}=(a_{n})_{n\geq 1}\in\Omega\setminus\overline{\mathsf{R}}$ a sequence $\mathbf{b}\in\overline{\mathsf{R}}$ for which $\overline{\Lambda}(\mathbf{b})=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}$.
Let $\sigma\colon\mathbb{Z}[i]^{\mathbb{N}}\to\mathbb{Z}[i]^{\mathbb{N}}$ be the left-shift map; that is, $\sigma(a_{1},a_{2},a_{3},\ldots)=(a_{2},a_{3},\ldots)$ for all $(a_{n})_{n\geq 1}\in\mathbb{Z}[i]^{\mathbb{N}}$. In [38], Yuri proved that $(\overline{\mathsf{R}},\sigma)$ is a sofic shift realized by a directed graph whose vertices are open prototype sets (see definition in Section 3) and with edges labeled by Gaussian integers. Yuri’s paper concerns the ergodic theory of a class of dynamical systems defined in compact subsets of $\mathbb{R}^{n}$, $n\in\mathbb{N}$. She considered measures equivalent to the $n$-dimensional Lebesgue measure and, thus, certain Lebesgue-null sets and some subsets of the associated shift space are discarded. These omissions force us to make a more detailed study of the relationship between $(\overline{\mathsf{R}},\sigma)$ and the dynamics of Hurwitz continued fraction when dealing with other problems such as those in Section 4 and Section 6.
@@
In fact, we will establish a more robust result by categorizing the set of normal numbers within the framework of the Borel hierarchy. The Borel hierarchy serves as a natural rank, quantifying the number of quantifiers required to construct a Borel set, commencing from open sets. In some sense, this measures the difficulty of determining whether a complex number is a Hurwitz normal number. A related result was obtained by Dani and Nogueira when they were studying the image of $\mathbb{Z}[i]^{2}$ under binary quadratic forms [13, Propositions 8.5, 8.6].
@@
The subshift $(\overline{\mathsf{R}},\sigma)$ is not a full shift, in other words, $\overline{\mathsf{R}}$ is not of the form $\mathscr{A}^{\mathbb{N}}$ for some $\mathscr{A}\subseteq\mathbb{Z}[i]$. Nonetheless, Theorem 1.3 shows that the system $(\overline{\mathsf{R}},\sigma)$ still has a considerable level of freedom in terms of specification. Bowen introduced the specification property in his seminal paper [6]. Bowen’s Specification Theorem asserts that all transitive hyperbolic systems possess the “specification property”, wherein complete independence is established among finite subsets with bounded gaps. The specification and its variations have proven to be potent tools for unlocking access to other properties, such as the uniqueness of measures of maximal entropy [6, 11], universality [32], and the intricate structure of generic points [4, 14].
The second application of Theorem 1.1 is related to transcendental numbers. In [2], Adamczewski and Bugeaud solved Cobham’s Conjecture when they proved that, for any $b\in\mathbb{N}_{\geq 2}$, a real number whose expansion in base $b$ is non-periodic and generated by finite automaton must be transcendental. They extended their work to continued fractions [1] and, afterward, in [8], Bugeaud gave a combinatorial condition on non-periodic continued fraction ensuring its transcendence. This condition can be expressed extremely neatly using the repetition exponent introduced by Bugeaud and Kim [10]. Given a finite alphabet $\mathscr{A}$ and a sequence $\mathbf{a}=(a_{n})_{n\geq 1}$ in $\mathscr{A}$, for each $n\in\mathbb{N}$ define
@@
Bugeaud proved that for any bounded non-periodic sequence $\mathbf{a}=(a_{n})_{n\geq 1}\in\mathbb{N}^{\mathbb{N}}$, if $\operatorname{rep}(\mathbf{a})<\infty$, the continued fraction (1) is transcendental for any $a_{0}\in\mathbb{Z}$ [8, Theorem 1.3]. In [19], the second author obtained a complex version of this result under the addition assumption $\min_{n}|a_{n}|\geq\sqrt{8}$. In this paper, we apply the techniques of the proof of Theorem 1.1 to construct a family of transcendental numbers whose partial quotients $(a_{n})_{n\geq 1}$ are bounded and such that $|a_{n}|=2$ for infinitely many $n\in\mathbb{N}$.
@@
The structure of the paper is as follows. First, in Section 2, we establish notation and conventions. In Section 3, we introduce the basic definitions of Hurwitz continued fractions, we discuss the classification of cylinders and prototype sets, and we give some basic results. In Section 4 we prove Theorem 1.1 and state without proof Proposition 7.1, our main technical result. Section 5 starts with some basic notions on Borel Hierarchy and the feeble-specification property. Later, we show Theorem 1.3 and then Theorem 5.1 (which implies Theorem 1.2). In Section 6 we show Theorem 1.4. Lastly, Section 7 contains the proof of Proposition 7.1.
@@ 2. Notation and conventions
    Let $\mathsf{rot},\mathsf{mir}_{1}:\mathbb{C}\to\mathbb{C}$ be given by $\mathsf{rot}(z)=iz$, $\mathsf{mir}_{1}(z)=\overline{z}$ and let $\mathsf{Di}_{8}$ be the group of isometries of $\mathbb{C}$ generated by $\mathsf{rot}$ and $\mathsf{mir}_{1}$. Put $\mathsf{mir}_{2}=\mathsf{rot}\circ\mathsf{rot}\circ\mathsf{mir}_{1}$ and denote by $\mathsf{Mir}$ the subgroup of $\mathsf{Di}_{8}$ generated by $\mathsf{mir}_{1}$ and $\mathsf{mir}_{2}$.
@@
    Given finitely many functions $h_{1},\ldots,h_{n}$ that are either $\iota$, $\tau_{a}$ for some $a\in\mathbb{Z}[i]$ or some element in $\mathsf{Di}_{8}$, we write $h_{n}\cdots h_{1}:=h_{n}\circ\cdots\circ h_{1}$.
  8. viii.
@@
    If $X$ is a topological space and $A\subseteq X$, we let $\operatorname{Cl}_{X}(A)$ (resp. $\operatorname{Int}_{X}(A)$ ) denote the closure of $A$ (resp. the interior). We omit the sub-index $X$ if the space is clear.
@@
    The term word typically refers to a finite set of ordered symbols, while a sequence denotes an infinite one. However, we use both terms interchangeably in certain situations where the set may be finite or infinite.
  11. xi.
@@
We implicitly use formulas for the inversion of discs and lines in the complex plane throughout the paper. Namely, for $k\in\mathbb{R}$, we have
@@ 3. Hurwitz continued fractions
$\forall z\in\mathbb{C}\quad T(z)=\begin{cases}z^{-1}-[z^{-1}],&\text{ if }z\neq 0,\\ 0,&\text{ if }z=0.\end{cases}$
For any $z\in\mathfrak{F}\setminus\{0\}$, we define $a_{0}(z)\colon=0$, $a_{1}(z)\colon=[z^{-1}]$ and, if $T^{n-1}(z)\neq 0$, we put $a_{n}(z)\colon=a_{1}(T^{n-1}(z))$ for $n\in\mathbb{N}$. As it is customary, the exponent in $T$ denotes iteration. For a general $z\in\mathbb{C}$, we define $a_{0}(z)\colon=[z]$ and $a_{n}(z)\colon=a_{n}(z-[z])$, $n\in\mathbb{N}$. The Hurwitz continued fraction (HCF) of $z$ is
(4) $[a_{0}(z);a_{1}(z),a_{2}(z),\ldots]_{\mathbb{C}}\colon=a_{0}(z)+\cfrac{1}{a_{1}(z)+\cfrac{1}{a_{2}(z)+\cfrac{1}{\ddots}}}.$
It was shown by Hurwitz [21] that $(a_{n}(z))_{n\geq 0}$ is infinite if and only if $z\in\mathbb{C}\setminus\mathbb{Q}(i)$. In this case, equation (4) can be interpreted as a limit. In the following, we restrict our attention to $z\in\mathfrak{F}$. For such $z$
@@
For such $\xi$, we have defined $[\xi]$ as the solution of (8) with the largest real and imaginary parts. In many situations, the breaking ties criterion is irrelevant for it applies on a set of Hausdorff dimension $1$, hence Lebesgue null. However, it ensures the uniqueness of the HCF expansion. If we are willing to loosen the expansion’s uniqueness slightly, Theorem 1.1 and its proof imply that it makes no difference what solution of (8) we choose.
We will use continued fractions which are not necessarily HCFs. We denote them by
@@
We reserve the notation $[b_{0};b_{1},b_{2},\ldots]_{\mathbb{C}}$ for HCFs.
@@
Take $z=[0;a_{1},a_{2},\ldots]_{\mathbb{C}}\in\mathfrak{F}\setminus\mathbb{Q}(i)$ and $(p_{n})_{n\geq 0}$, $(q_{n})_{n\geq 0}$ as in (6).
  1. (i).
@@
  4. (iv).
    [29, Theorem 1] We have
@@
    $\forall n\in\mathbb{N}\quad\left|z-\frac{p_{n}}{q_{n}}\right|<\frac{\lambda}{|q_{n}q_{n+1}|}.$
  6. (vi).